# Semi-empirical Modeling of Supermassive Black Hole Evolution: Highlighting a possible tension between Demographics and Gravitational Wave Background

### v1

@@ 3 Data and analysis
$\epsilon_{\rm thin}$ $\mathcal{U}(0.06,0.42)$ 2.4
$\log\lambda_{0}$ $\mathcal{U}[-3,1]$ 2.3
$\xi$ $\mathcal{U}[0,3]$ 2.3
$\eta$ $\mathcal{U}[-1,1]$ 2.11
$\log f_{\bullet\bullet\rightarrow\bullet}$ $\mathcal{U}[-3,1]$ or $\mathcal{N}(0,0.1)$ 2.13
$\sigma_{\log M_{\rm H}}$ $\mathcal{U}(0,1]$ 2.15
@@
Parameter $M_{\bullet}-M_{\star}$ [obs.] $M_{\bullet}-\sigma$ [obs.] $M_{\bullet}-\sigma$ [deb.] $M_{\bullet}-M_{\star}$ [obs.] $+f_{\bullet\bullet\rightarrow\bullet}\approx 1$
[obs.] [obs.] [deb.] $+f_{\bullet\bullet\rightarrow\bullet}\approx 1$
$\epsilon_{\rm thin}$ $0.066^{+0.003}_{-0.008}$ $0.11^{+0.01}_{-0.01}$ $0.34^{+0.02}_{-0.03}$ $0.066^{+0.003}_{-0.008}$
$\log\lambda_{0}$ $-2.25^{+0.26}_{-0.23}$ $-1.43^{+0.22}_{-0.15}$ $-0.62^{+0.12}_{-0.08}$ $-2.22^{+0.27}_{-0.22}$
@@
$\log f_{\bullet\bullet\rightarrow\bullet}$ $-0.77^{+0.8}_{-0.5}$ $-0.85^{+0.9}_{-0.6}$ $-1.4^{+1.1}_{-0.9}$ $-0.04^{+0.09}_{-0.09}$
@@
Table 3: Results of our Bayesian analysis in terms of marginalized posterior estimates. For every parameter (listed in the first column) of our semi-empirical model we report the median and the $2\sigma$ credible interval. The different columns refer to the different determinations of the local BH mass functions discussed in the main text. The last line lists the reduced $\chi^{2}_{r}$ of the overall fit in the various cases.
@@ 3.1.1 Parameter estimates
The radiative efficiency $\epsilon_{\rm thin}$ is a well determined parameter, though it is very sensitive to the adopted local BH mass function determination: values $\epsilon_{\rm thin}\approx 0.06$ close to the lower bound allowed for a thin disk are required when adopting the BH mass function based on the observed $M_{\bullet}-M_{\star}$ relationship; values $\epsilon_{\rm thin}\approx 0.1$ around the usually adopted one for the average population of AGNs (e.g., [59, 61]) are found when basing on the observed $M_{\bullet}-\sigma$ relationship; extreme values $\epsilon_{\rm thin}\approx 0.3$ close to the upper bound allowed for a thin disk are required when adopting the BH mass function based on the debiased $M_{\bullet}-\sigma$ relationship. These variations are easily understood basing on the classic Soltan [57, 58] argument. As it is more easily seen in the simple case of Equation (2.7), the local BH mass density $\rho_{\bullet}\equiv\int{\rm d}\log M_{\bullet}\,M_{\bullet}\,N(\log M_{\bullet},t)$ is basically determined by the integral over cosmic time of the AGN luminosity density $\rho_{\bullet}\propto\frac{1-\epsilon-\epsilon_{\rm kin}}{\epsilon}\,\int{\rm d}t\,\int{\rm d}\log L_{\rm AGN}\,L_{\rm AGN}\,N(\log L_{\rm AGN},t)$ scaled by the inverse of the radiative efficiency $\epsilon^{-1}$. Thus BH mass function determinations featuring a higher normalization at the knee, which implies a larger BH mass density $\rho_{\bullet}$, require lower average efficiencies.
@@
The values of the average efficiency estimated here could have far reaching physical consequences, since in the standard picture of disk accretion these can be traced back to the spinning state of the supermassive BHs [186, 187]. For example, if the local BH mass function from the debiased $M_{\bullet}-\sigma$ relationship would be the true one, the related high efficiencies $\epsilon\sim 0.3$ would correspondingly imply rapidly spinning behaviors in the typical supermassive BH population. Unfortunately, this argument could not be considered as conclusive given the present uncertainties inherent to the local mass function determination.
@@
The parameters $\lambda_{0}$ and $\xi$ determining the local normalization and redshift evolution of the mean Eddington ratio are also rather well determined by our analysis. Here the relevant point is the mapping between the AGN luminosity function and the BH mass function via the Eddington ratio distribution that enters the continuity equation approach. Specifically, lower average values of $\lambda$ tend to originate a flatter BH mass function at the high-mass end, since large BH masses correspond to moderate AGN luminosities (approximately $L_{\rm AGN}\propto\lambda\,M_{\bullet}$ holds) falling in the flatter portion of the AGN luminosity function. Thus local BH mass function determinations with a flatter (steeper) high mass end prefer lower (higher) values of the average $\lambda_{0}$. This explain the rising trend of the $\lambda_{0}$ estimate from the BH mass function determination based on the $M_{\bullet}-M_{\star}$ relationship, which is the flatter, toward that based on the debiased $M_{\bullet}-M_{\star}$ relationship, which feature the steeper shape.
@@ 3.1.2 BH demographics and clustering
In Figure 2 we illustrate the BH mass function resulting from our semi-empirical model. Solid colored lines show the bestfits to the three local BH mass function determinations (colored symbols) based on the observed $M_{\bullet}-M_{\star}$ (red), observed $M_{\bullet}-\sigma$ (blue) and debiased $M_{\bullet}-\sigma$ (green) relationship. In addition, the magenta line is the bestfit to the $M_{\bullet}-M_{\star}$ determination when a narrow prior on the scaling parameter of the BH to galaxy merger rate $f_{\bullet\bullet-\rightarrow\bullet}$ around unity is adopted. All the fits are remarkably good, highlighting that the parameterizations adopted in our semi-empirical model are sensible and sufficient to describe the BH demographics. Moreover, the dashed colored lines illustrate the corresponding active BH mass function at high redshift $z\sim 5$, which should be compared with the observational estimates by [89, 90, 177, 88, 91, 68] at $z\sim 4-6$ (dark grey symbols). Note that the high-redshift active BH mass function has not been included in the fitting process, so the good agreement found with these independent data is a genuine prediction of our model, and further corroborates our choice of a redshift dependent mean Eddington ratio.
@@
In Figure 3 we compare the large-scale bias as a function of BH mass from our semi-empirical model for the different local BH mass function determinations (colored lines) to clustering data at $z\sim 0.01-0.25$ by [173, 174, 157, 158], finding an excellent agreement. In Figure 4 the related local BH to halo mass relationship is displayed, and compared with various observational estimates by [178, 157, 180, 179] derived by a variety of methods (e.g., clustering, kinematics, weak lensing). Actually the data are rather dispersed, partly because of intrinsic variance in the relationships, and partly because of possible observational systematics. In particular, there is a tendency for weak lensing estimates to predict a BH to halo mass relation with a lower normalization with respect to clustering measurements. Note that in view of these observational uncertainties, the BH to halo mass relation has not been fitted upon in our analysis. However, the relation derived via abundance matching from our semi-empirical model, also in view of the estimated intrinsic scatter $\sigma_{\log M_{\rm H}}\gtrsim 0.3-0.5$ dex (colored errorbars in the plot), is remarkably consistent with the cloud of datapoints.
@@
Figure 5 illustrates the local BH duty cycle as a function of BH mass from our semi-empirical model for the different local BH mass function determinations (colored lines), compared with classic observational estimates by [175, 176, 138] derived from AGN active fractions in the local Universe. First of all, it is important to notice that there is a rather large dispersion in the data, which partly reflects differences in bolometric corrections, luminosity thresholds, and systematic offsets in the BH mass estimates; for these reasons the duty cycle has not been fitted upon in our analysis. Nevertheless, the duty cycle reconstructed from the continuity equation is found to be broadly consistent with these observational estimates. More in detail, higher duty cycle around $1-10\%$ are produced for the models fitting the local BH mass function determination based on the observed $M_{\bullet}-M_{\star}$ relation, while smaller duty cycles $\sim 0.1-1\%$ are originated in the model fitting BH mass function based on the debiased $M_{\bullet}-\sigma$ relationship.
@@
In Figure 6 we showcase the mean Eddington ratio as a function of redshift as derived from our analysis for the different local BH mass function determinations (colored lines). These outcomes are compared with a multitude of observational estimates from X-ray/mid-IR/optical AGNs by [65, 181, 182, 76], broad-absorption-line quasars by [89, 90, 37, 39], and little red dots by [84, 183, 88, 184]. The various datasets are extremely heterogeneous, so that the comparison with the model outcomes should be taken only as indicative. In fact, we stress that the parameters $\lambda_{0}$ and $\xi$ determining in our model the redshift-dependent Eddington ratio after Equation (2.3) have not been fitted on the data reported here, but instead have been determined by reproducing the BH demographics and clustering via the continuity equation approach. Nevertheless, it is reassuring that the model outputs on $\lambda(z)$ are broadly consistent with these observational determinations for individual AGNs. As already mentioned, the outcome referring to the BH mass function determination based on the observed $M_{\bullet}-M_{\star}$ relation tends to predict rather low mean Eddington ratios $\lambda\lesssim 0.01$ in the local Universe and a steep rise toward high redshift values $\lambda\sim 1$ at $z\sim 6$. On the other hand, the model outcomes referring to the BH mass function determinations based on the observed or debiased $M_{\bullet}-\sigma$ relations point toward higher current normalization $\lambda\lesssim 0.3$ and an appreciably weaker evolution toward high redshift. This behavior reflects the well-known degeneracy between the local duty cycle and the evolution of the Eddington ratio (see [23]).
@@
In Figure 7 we illustrate the relative contributions of mergers to the increase in the BH mass function, sliced in different bins of BH mass and redshift. The color scale illustrates the ratio between the production rate of the mass function due to mergers (i.e., the positive term in the Smoluchowski equation) normalized to the sum of this quantity and of the analogous one due to accretion. It is evident that the main impact of mergers in increasing the mass function occurs toward low redshift $z\lesssim 1.5$ and for relatively high masses $M_{\bullet}\gtrsim$ some $10^{8}\,M_{\odot}$. Such a contribution is extended somewhat toward higher redshift and smaller BH masses for the BH mass function determination based on the $M_{\bullet}-M_{\star}$ relation, which features the highest normalization at the high mass end, and thus leave more rooms for redistribution of the masses there. An additional yet modest contribution occurs at intermediate redshift $z\sim 2-4$ because the redshift increase in the BH (and galaxy) merger rate cooperates with a still appreciably high BH mass function; at higher $z$ the effect is less evident since the merger rate stays put but the BH mass function drops considerably. Finally, it is worth noticing that for very large BH masses $M_{\bullet}\gtrsim 10^{9}\,M_{\odot}$ mergers are quite relevant even at substantial redshifts $z\gtrsim 6$, though clearly such over-massive BHs in the primordial Universe would be quite rare.
3.1.3 GW stochastic background
@@
In Figure 8 we illustrate our posterior estimates for the normalized stretch amplitude of the nano-Hz GW background for the different local BH mass function determinations (colored violin plots), with the black dots and solid errorbars illustrating the median and the $2\sigma$ credible intervals. These should be compared with the observational estimate of the background amplitude by the NANOGrav collaboration [96] (yellow solid line with shaded area). It appears evident that the outcomes of our analysis do not favor a substantial contribution to the observed GW background from supermassive BH mergers. Even in the extreme case (magenta violin plot) where we have adopted the BH determination based on the $M_{\bullet}-M_{\star}$ relation (which has the highest normalization) and we have imposed a narrow prior around unity on the rescaling parameter $f_{\bullet\bullet\rightarrow\bullet}$ between the BH and the galaxy merger rate, our analysis predicts that only a fraction $\lesssim 30-50\%$ (at $\sim 3\sigma$) of the observed background can be traced back to supermassive BH mergers. Our results are in good agreement with the outcomes of the Astrid and TNG-300 numerical simulations (see [112]) and of the holodeck suite (see [96]); in fact, the latter two have a median closer to our maximized result with $f_{\bullet\bullet\rightarrow\bullet}\approx 1$, but the large uncertainties make them consistent also with our more conservative estimates.
@@
In Figure 9 we highlight the contribution of supermassive BH mergers to the nano-Hz GW stochastic background sliced in redshift and BH mass, for the different local BH mass function determinations (different panels). Most of background is contributed by relatively low redshift $z\lesssim 3$ sources and large BH masses $M_{\bullet}\gtrsim 10^{9}\,M_{\odot}$. Such a contribution extends toward higher redshift and smaller BH masses for the BH mass function determination based on the $M_{\bullet}-M_{\star}$ relation, which features the highest normalization at the high mass end; this behavior is then reinforced when the binary BH merger rate is maximized and imposed to follow the galaxy merger rate (i.e., when $f_{\bullet\bullet\rightarrow\bullet}\approx 1$; top right panel). This figure also highlights that the tension between BH demographics and GW background is robust to residual uncertainties on the basic inputs of our data-driven model. For example, one could implement variations in the AGN luminosity function at $z\gtrsim 4.5$ (see Appendix A), which may account for a steeper trend/higher normalization with respect to [74] as claimed by [188, 189, 190, 191] or possibly suggested by an increased UV space density due to little red dots in case these are truly AGN dominated (see [91]). However, the impact on the GW background from such changes at relatively high redshift and small to moderate AGN luminosities/BH masses would be minor (more quantitative analysis is presented in Appendix B).
@@
A possible tension between the GW background expected from BH mergers and the ones measured by PTA has been noted and discussed before in the literature (e.g., [103, 105, 96, 154, 192, 108, 193, 112]). However, these studies lack a self-consistent treatment of BH mergers, in that the related rates (usually derived from galaxy merger rates via a BH-galaxy scaling relation) are used to compute the expected GW background, but are not employed at the same time to build up and check the related impact on the mass function. Our analysis point out the tension via a semi-empirical, data-driven model based on the continuity plus Smoluchowski equation, that treats unitarily accretion and mergers, incorporates the latest observational determinations of the AGN luminosity functions and galaxy merger rates, and is jointly constrained by local BH demographics and GW background.
@@
The situation is better clarified in Figure 10, where we illustrate the merger rate per descendant BH (basically Equation 2.11 integrated over the mass ratio distribution) vs. redshift, referred to a reference BH mass $M_{\bullet}\sim 10^{10}\,M_{\odot}$ which is taken as a representative contributor to the GW background. Note that the mass dependence is anyway not an issue since it is weak and the range of BH masses contributing to the background is rather limited (see above). Colored solid lines illustrate the median outcomes from our semi-empirical model and MCMC analysis, for the different local BH mass function determinations. Essentially they are rescaled version (by the parameter $f_{\bullet\bullet\rightarrow\bullet}$) of the solid gray curve, which represents the galaxy merging rate as estimated from pairs counts by JWST. However, note that the $2\sigma$ uncertainty (not reported for the sake of clarity otherwise the plot would look crowded) for the red, blue and green cases essentially extends from the magenta line downwards, since our model can only set an upper limit to the merging rates per BH, as it is evident from the posterior of the parameter $f_{\bullet\bullet\rightarrow\bullet}$ in Fig. 1. For comparison, the dashed gray line refers to the major merger rate per galaxy from the Illustris simulation by [194], converted to the reference BH mass via the BH-stellar mass relation by [47], while the dotted gray line is the major merger rate per dark matter halo by [195], converted to the reference BH mass via the BH-halo mass relation by [162].
@@
The take-home message is that to reproduce the high mass end of the estimated local mass functions (regardless of the specific determination), our model restricts the merger rate to be smaller or in any case not to exceed that inferred from galaxy pairs by [152, 153]. This is in turn slightly lower at $z\lesssim 0.5$ than that measured for galaxies in hydrodynamic simulations and for halos in $N-$body codes. Contrariwise, fitting the GW background measured by PTA experiments would require the merger rate per BH to be appreciably higher, as illustrated by the yellow curves that refer to various parametric analyses of the PTA data by [192, 154, 193]. However, these high rates would inevitably corrupt the local BH mass function at the high mass end, in a way that cannot be effectively cured by any reasonable accretion-related parameter set. Moreover, these merger rates per BH required to fit the PTA data turn out to be appreciably higher than the merger rate per galaxies or per halo expected from observations of galaxy pairs and from numerical simulations, an occurrence which would be in itself rather problematic to explain.

### v2

@@ 3 Data and analysis
$\log\lambda_{0}$ $\mathcal{U}[-3,1]$ 2.5
$\xi$ $\mathcal{U}[0,3]$ 2.5
$\epsilon_{\rm thin}$ $\mathcal{U}(0.06,0.42)$ 2.7
$\eta$ $\mathcal{U}[-1,1]$ 2.13
$\log f_{\bullet\bullet\rightarrow\bullet}$ $\mathcal{U}[-3,1]$ or $\mathcal{N}(0,0.1)$ 2.15
$\sigma_{\log M_{\rm H}}$ $\mathcal{U}(0,1]$ 2.17
@@
Parameter $M_{\bullet}-M_{\star}$ [obs.] $M_{\bullet}-\sigma$ [obs.] $M_{\bullet}-\sigma$ [deb.] $M_{\bullet}-M_{\star}$ [obs.] $f_{\bullet\bullet\rightarrow\bullet}\approx 1$
$\epsilon_{\rm thin}$ $0.066^{+0.003}_{-0.008}$ $0.11^{+0.01}_{-0.01}$ $34^{+0.02}_{-0.03} 0.0$$066^{+0.003}_{-0.008}$
$\log\lambda_{0}$ $2.25^{+0.26}_{-0.23}$ $1.43^{+0.22}_{-0.15}$ $0.62^{+0.12}_{-0.08}$ $2.22^{+0.27}_{-0.22}$
@@
$\log f_{\bullet\bullet\rightarrow\bullet}$ $0.77^{+0.8}_{-0.5}$ $0.85^{+0.9}_{-0.6}$ $1.4^{+1.1}_{-0.9}$ $0.04^{+0.09}_{-0.09}$
@@
Table 3: Results of our Bayesian analysis in terms of marginalized posterior estimates. For every parameter (listed in the first column) of our semi-empirical framework we report the median and the $2\sigma$ credible interval. The different columns refer to the different determinations of the local BH mass functions discussed in the main text. The last line lists the reduced $\chi^{2}_{r}$ of the overall fit in the various cases.
@@ 3.1.1 Parameter estimates
The radiative efficiency $\epsilon_{\rm thin}$ is a well determined parameter, though it is very sensitive to the adopted local BH mass function determination: values $\epsilon_{\rm thin}\approx 0.06$ close to the lower bound allowed for a thin disk are required when adopting the BH mass function based on the observed $M_{\bullet}-M_{\star}$ relationship; values $\epsilon_{\rm thin}\approx 0.1$ around the usually adopted one for the average population of AGNs (e.g., [Davis2011, Trakhtenbrot2017]) are found when basing on the observed $M_{\bullet}-\sigma$ relationship; extreme values $\epsilon_{\rm thin}\approx 0.3$ close to the upper bound allowed for a thin disk are required when adopting the BH mass function based on the debiased $M_{\bullet}-\sigma$ relationship. These variations are easily understood basing on the classic Soltan [Soltan1982, Salucci1999] argument. As it is more easily seen in the simple case of Equation (2.9), the local BH mass density $\rho_{\bullet}\equiv\int{\rm d}\log M_{\bullet}\,M_{\bullet}\,N(\log M_{\bullet},t)$ is essentially determined by the integral over cosmic time of the AGN luminosity density $\rho_{\bullet}\propto\frac{1-\epsilon-\epsilon_{\rm kin}}{\epsilon}\,\int{\rm d}t\,\int{\rm d}\log L_{\rm AGN}\,L_{\rm AGN}\,N(\log L_{\rm AGN},t)$ scaled by the inverse of the radiative efficiency $\epsilon^{-1}$. Thus BH mass function determinations featuring a higher normalization at the knee, which implies a larger BH mass density $\rho_{\bullet}$, require lower average efficiencies.
@@
The values of the average efficiency estimated here could have far reaching physical consequences, since in the standard picture of disk accretion these can be traced back to the spinning state of the supermassive BHs [Bardeen1972, Zhang2019]. For example, if the local BH mass function from the debiased $M_{\bullet}-\sigma$ relationship would be the true one, the related high efficiencies $\epsilon\sim 0.3$ would correspondingly imply rapidly spinning behaviors in the typical supermassive BH population. Unfortunately, this argument could not be considered as conclusive given the present uncertainties inherent to the local mass function determination.
@@
The parameters $\lambda_{0}$ and $\xi$ determining the local normalization and redshift evolution of the mean Eddington ratio are also rather well determined by our analysis. Here the relevant point is the mapping between the AGN luminosity function and the BH mass function via the Eddington ratio distribution that enters the continuity equation approach. Specifically, lower average values of $\lambda$ tend to originate a flatter BH mass function at the high-mass end, since large BH masses correspond to moderate AGN luminosities (approximately $L_{\rm AGN}\propto\lambda\,M_{\bullet}$ holds) falling in the flatter portion of the AGN luminosity function. Thus local BH mass function determinations with a flatter (steeper) high mass end prefer lower (higher) values of the average $\lambda_{0}$. This explain the rising trend of the $\lambda_{0}$ estimate from the BH mass function determination based on the $M_{\bullet}-M_{\star}$ relationship, which is the flatter, toward that based on the debiased $M_{\bullet}-M_{\star}$ relationship, which features the steeper shape.
@@ 3.1.2 BH demographics and clustering
In Figure 4 we illustrate the BH mass function resulting from our semi-empirical framework. Solid colored lines show the bestfits to the three local BH mass function determinations (colored symbols) based on the observed $M_{\bullet}-M_{\star}$ (red), observed $M_{\bullet}-\sigma$ (blue) and debiased $M_{\bullet}-\sigma$ (green) relationship. In addition, the magenta line is the bestfit to the $M_{\bullet}-M_{\star}$ determination when a narrow prior on the scaling parameter of the BH to galaxy merger rate $f_{\bullet\bullet-\rightarrow\bullet}$ around unity is adopted. All the fits are remarkably good, highlighting that the parameterizations adopted in our semi-empirical framework are sensible and sufficient to describe the BH demographics. Moreover, the dashed colored lines illustrate the corresponding active BH mass function at high redshift $z\sim 5$, which should be compared with the observational estimates by [He2024, Lai2024, Taylor2025, Matthee2024, Kokorev2024, Wu2022] at $z\sim 4-6$ (dark grey symbols). Note that the high-redshift active BH mass function has not been included in the fitting process, so the good agreement found with these independent data is a genuine prediction of our model, and further corroborates our choice of a redshift dependent mean Eddington ratio.
In Figure 5 we compare the large-scale bias $b(M_{\bullet},z)$ as a function of BH mass from our semi-empirical framework for the different local BH mass function determinations (colored lines) to clustering data at $z\sim 0.01-0.25$ by [Powell2018, Allevato2021, Krumpe2015, Krumpe2023], finding an excellent agreement. In addition, the inset shows the BH bias as a function of redshift, after averaging over the supermassive BH mass function, i.e. computing $\langle b\rangle(z)\equiv\int{\rm d}\log M_{\bullet}\,N(\log M_{\bullet},z)\,b(M_{\bullet},z)/$ $\int{\rm d}\log M_{\bullet}\,N(\log M_{\bullet},z)$. Although this redshift dependence of the bias has not been fitted upon, the outcomes from our semi-empirical framework agree reasonably well with the observational estimates from optical (plus signs; [Shen2009, White2012, Ross2020]), X-ray (diamonds; [Allevato2011, Allevato2014, Allevato2016]) and radio-selected (inverse triangles; [Magliocchetti2017, Mazumder2022, Hale2025]) samples.
@@
In Figure 6 the related local BH to halo mass relationship is displayed, and compared with various observational estimates by [Marasco2021, Krumpe2015, Li2024, Robinson2021] derived by a variety of methods (e.g., clustering, kinematics, weak lensing). Actually the data are rather dispersed, partly because of intrinsic variance in the relationships, and partly because of possible observational systematics. In particular, there is a tendency for weak lensing estimates to predict a BH to halo mass relation with a lower normalization with respect to clustering measurements. Note that in view of these observational uncertainties, the BH to halo mass relation has not been fitted upon in our analysis. However, the relation derived via abundance matching from our semi-empirical framework, also in view of the estimated intrinsic scatter $\sigma_{\log M_{\rm H}}\gtrsim 0.3-0.5$ dex (colored errorbars in the plot), is consistent with the cloud of datapoints.
@@
Figure 7 illustrates the local BH duty cycle as a function of BH mass from our semi-empirical framework for the different local BH mass function determinations (colored lines), compared with classic observational estimates by [Kauffmann2003, Goulding2010, Schulze2010] derived from AGN active fractions in the local Universe. First of all, it is important to notice that there is a rather large dispersion in the data, which partly reflects differences in bolometric corrections, luminosity thresholds, and systematic offsets in the BH mass estimates; for these reasons the duty cycle has not been fitted upon in our analysis. Nevertheless, the duty cycle reconstructed from the continuity equation is found to be broadly consistent with these observational estimates. More in detail, higher duty cycle around $1-10\%$ are produced for the models fitting the local BH mass function determination based on the observed $M_{\bullet}-M_{\star}$ relation, while smaller duty cycles $\sim 0.1-1\%$ are originated in the model fitting BH mass function based on the debiased $M_{\bullet}-\sigma$ relationship.
@@
In the top panel of Figure 8 we showcase the mean Eddington ratio as a function of redshift as derived from our analysis for the different local BH mass function determinations (colored lines). These outcomes are compared with a multitude of observational estimates from X-ray/mid-IR/optical AGNs by [Vestergaard2009, Nobuta2012, Dai2014, Duras2020], broad-absorption-line quasars by [He2024, Lai2024, Farina2022, Fan2023], and little red dots by [Harikane2023, Greene2024, Matthee2024, Maiolino2024]. The various datasets are somewhat heterogeneous, so that the comparison with the model outcomes should be taken only as indicative. In fact, we stress that the parameters $\lambda_{0}$ and $\xi$ determining in our model the redshift-dependent Eddington ratio after Equation (2.5) have not been fitted on the data reported here, but instead have been determined by reproducing the BH demographics and clustering via the continuity equation approach. Nevertheless, it is reassuring that the model outputs on $\lambda(z)$ are broadly consistent with these observational determinations for individual AGNs. As already mentioned, the outcome referring to the BH mass function determination based on the observed $M_{\bullet}-M_{\star}$ relation tends to predict rather low mean Eddington ratios $\lambda\lesssim 0.01$ in the local Universe and a steep rise toward high redshift values $\lambda\sim 1$ at $z\sim 6$. This behavior agrees well with the average Eddington ratio observationally determined for type-1 AGNs up to $z\lesssim 4$ by [Kelly2013], as fitted by [Tucci2017]; it also accords with the dependence assumed on phenomenological grounds by [Shankar2013]. On the other hand, the model outcomes referring to the BH mass function determinations based on the observed or debiased $M_{\bullet}-\sigma$ relations point toward somewhat higher current normalization $\lambda\lesssim 0.3$ and an appreciably weaker evolution toward high redshift. This behavior reflects the well-known degeneracy between the local duty cycle and the evolution of the Eddington ratio (see [Shankar2013]). In the bottom panel of Figure 8 we illustrate the full Eddington ratio distribution at three representative redshifts $z\approx 0.8$ (left) $1.8$ (middle) and $3.2$ (right). The outcomes of our semi-empirical framework are in good accord with the estimates for type-1 AGNs by [Kelly2013] (circles) and with the corresponding lognormal fits by [Tucci2017] (dashed lines).
@@
In Figure 9 we illustrate the relative contributions of mergers to the increase in the BH mass function, sliced in different bins of BH mass and redshift. The color scale illustrates the ratio between the production rate of the mass function due to mergers (i.e., the positive term in the Smoluchowski equation) normalized to the sum of this quantity and of the analogous one due to accretion. It is evident that the main impact of mergers in increasing the mass function occurs toward low redshift $z\lesssim 1.5$ and for relatively high masses $M_{\bullet}\gtrsim$ some $10^{8}\,M_{\odot}$. Such a contribution is extended somewhat toward higher redshift and smaller BH masses for the BH mass function determination based on the $M_{\bullet}-M_{\star}$ relation, which features the highest normalization at the high mass end, and thus leave more rooms for redistribution of the masses there. An additional yet modest contribution occurs at intermediate redshift $z\sim 2-4$ because the redshift increase in the BH (and galaxy) merger rate cooperates with a still appreciably high BH mass function; at higher $z$ the effect is less evident since the merger rate stays put but the BH mass function drops considerably. Finally, it is worth noticing that for very large BH masses $M_{\bullet}\gtrsim 10^{9}\,M_{\odot}$ mergers are quite relevant even at substantial redshifts $z\gtrsim 6$, though clearly such over-massive BHs in the primordial Universe would be rare.
3.1.3 Stochastic GW background
@@
In Figure 10 we illustrate our posterior estimates for the normalized stretch amplitude of the nano-Hz GW background for the different local BH mass function determinations (colored violin plots), with the black dots and solid errorbars illustrating the median and the $2\sigma$ credible intervals. These should be compared with the observational estimate of the background amplitude by the NANOGrav collaboration [Agazie2023] (yellow solid line with shaded area). It appears evident that the outcomes of our analysis do not favor a substantial contribution to the observed GW background from supermassive BH mergers. Even in the extreme case (magenta violin plot) where we have adopted the BH determination based on the $M_{\bullet}-M_{\star}$ relation (which has the highest normalization) and we have imposed a narrow prior around unity on the rescaling parameter $f_{\bullet\bullet\rightarrow\bullet}$ between the BH and the galaxy merger rate, our analysis predicts that only a fraction $\lesssim 30-50\%$ (at $\sim 3\sigma$) of the observed background can be traced back to supermassive BH mergers. Our results are in good agreement with the outcomes of the Astrid and TNG-300 numerical simulations (see [Chen2025]) and of the holodeck suite (see [Agazie2023]); in fact, the latter two have a median closer to our maximized result with $f_{\bullet\bullet\rightarrow\bullet}\approx 1$, but the large uncertainties make them consistent also with our more conservative estimates.
@@
In Figure 11 we highlight the contribution of supermassive BH mergers to the nano-Hz GW stochastic background sliced in redshift and BH mass, for the different local BH mass function determinations (different panels). Most of background is contributed by relatively low redshift $z\lesssim 3$ sources and large BH masses $M_{\bullet}\gtrsim 10^{9}\,M_{\odot}$. Such a contribution extends toward higher redshift and smaller BH masses for the BH mass function determination based on the $M_{\bullet}-M_{\star}$ relation, which features the highest normalization at the high mass end; this behavior is then reinforced when the binary BH merger rate is maximized and imposed to follow the galaxy merger rate (i.e., when $f_{\bullet\bullet\rightarrow\bullet}\approx 1$; top right panel). This figure also highlights that the tension between BH demographics and GW background is robust to residual uncertainties on the basic inputs of our data-driven model. For example, one could implement variations in the AGN luminosity function at $z\gtrsim 4.5$ (see Appendix A), which may account for a steeper trend/higher normalization with respect to [Shen2020] as claimed by [Giallongo2019, Kulkarni2019, Barlow2023, Barlow2025] or possibly suggested by an increased UV space density due to little red dots in case these are truly AGN dominated (see [Kokorev2024]). However, the impact on the GW background from such changes at relatively high redshift and small to moderate AGN luminosities/BH masses would be minor (more quantitative analysis is presented in Appendix B).
@@
A possible tension between the GW background expected from BH mergers and the one measured by PTA has been noted and discussed before in the literature (e.g., [Sesana2016, Chen2019, Agazie2023, Agazie2023b, Middleton2021, Villalba2022, SatoPolito2025, Casey2025, Chen2025, Kis2025]). However, these studies lack a self-consistent treatment of BH mergers, in that the related rates (usually derived from galaxy merger rates via a BH-galaxy scaling relation) are used to compute the expected GW background, but are not employed at the same time to build up and check the related impact on the mass function. Our analysis point out the tension via a semi-empirical, data-driven model based on the continuity plus Smoluchowski equation, that treats unitarily accretion and mergers, incorporates the latest observational determinations of the AGN luminosity functions and galaxy merger rates, and is jointly constrained by local BH demographics and GW background.
@@
The situation is better clarified in Figure 12, where we illustrate the merger rate per descendant BH (basically Equation 2.13 integrated over the mass ratio distribution) vs. redshift, referred to a reference BH mass $M_{\bullet}\sim 10^{10}\,M_{\odot}$ which is taken as a representative contributor to the GW background. Note that the mass dependence is anyway not an issue since it is weak and the range of BH masses contributing to the background is rather limited (see above). Colored solid lines illustrate the median outcomes from our semi-empirical framework and MCMC analysis, for the different local BH mass function determinations. Essentially they are rescaled version (by the parameter $f_{\bullet\bullet\rightarrow\bullet}$) of the solid gray curve, which represents the galaxy merging rate as estimated from pairs counts by JWST. However, note that the $2\sigma$ uncertainty (not reported for the sake of clarity otherwise the plot would look crowded) for the red, blue and green cases essentially extends from the magenta line downwards, since our model can only set an upper limit to the merging rates per BH, as it is evident from the posterior of the parameter $f_{\bullet\bullet\rightarrow\bullet}$ in Fig. 3. For comparison, the dashed gray line refers to the major merger rate per galaxy from the Illustris simulation by [Rodriguez2015], converted to the reference BH mass via the BH-stellar mass relation by [Kormendy2013], while the dotted gray line is the major merger rate per dark matter halo by [Fakhouri2010], converted to the reference BH mass via the BH-halo mass relation by [Ferrarese2002].
@@
The take-home message is that to reproduce the high mass end of the estimated local mass functions (regardless of the specific determination), our model restricts the merger rate to be smaller or in any case not to exceed that inferred from galaxy pairs by [Duan2025, Puskas2025]. This is in turn slightly lower at $z\lesssim 0.5$ than that measured for galaxies in hydrodynamic simulations and for halos in NN-body codes. Contrariwise, fitting the GW background measured by PTA experiments would require the merger rate per BH to be appreciably higher, as illustrated by the yellow curves that refer to various parametric analyses of the PTA data by [Middleton2021, Agazie2023b, Casey2025]. However, these high rates would inevitably corrupt the local BH mass function at the high mass end, in a way that cannot be effectively cured by any reasonable accretion-related parameter set. Moreover, these merger rates per BH required to fit the PTA data turn out to be appreciably higher than the merger rate per galaxies or per halo expected from observations of galaxy pairs and from numerical simulations, an occurrence which would be in itself rather problematic to explain.